The hydronium content of a popular soft drink is measured and found to be 5.63×10⁻³ M. What is the pH of this soft drink?
1
Expert's answer
2020-05-07T14:15:15-0400

Solution:

[H3O+] = 5.63×10-3 M

The pH is calculated using the expression:

pH = - log[H3O+]

pH = - log(5.63×10-3) = 2.2495 = 2.25

pH = 2.25


Answer: The pH of this soft drink is 2.25